OpenStudy (anonymous):

Compute the limit of the sequence (n^2+1)^1/2/3n + 5n as n approaches infinity?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

Sorry its 3n + 5

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty}\frac{ \sqrt{n^2+1} }{ 3n+5 }\\ \lim_{n \rightarrow \infty}\frac{ \sqrt{n^2+1} }{3n +5}\\ \lim_{n \rightarrow \infty}\frac{ \sqrt{n^2}\sqrt{1+\frac{1}{n^2}} }{3n+5 }\\ \lim_{n \rightarrow \infty}\frac{ |n|\sqrt{1+\frac{1}{n^2}} }{3n+5 }\] Since \(n>0\), you have \(|n|=n\): \[\lim_{n \rightarrow \infty}\frac{ n\sqrt{1+\frac{1}{n^2}} }{3n+5 }\\ \lim_{n \rightarrow \infty}\frac{ \sqrt{1+\frac{1}{n^2}} }{3+\frac{5}{n} }\\ \]